Directions: Solve the radical equation. Check all propsed solutions.
square root y^2-3y+18=y+2
The square root is over y^2-3y+18, thats it, not the y+2

Answer by fractalier(2101)

Okay first we square both sides...we get
sqrt( y^2-3y+18 ) = y + 2
y^2 - 3y + 18 = y^2 + 4y + 4
now collect terms and solve...we get
-3y + 18 = 4y + 4
14 = 7y
y = 2
you have to check these by plugging 2 in for y in the original...turns out okay...